Find T4, T7 and T9 for each of the following geometric progressions :3, 6,12,24


1
Expert's answer
2021-03-30T07:41:11-0400
"t_1=3, t_2=6, t_3=12, t_4=24"

Common ratio "r=\dfrac{t_2}{t_1}=\dfrac{t_3}{t_2}=\dfrac{t_4}{t_3}=2"


"t_4=24"

"t_7=t_1\cdot r^6=3\cdot 2^6=192"

"t_9=t_1\cdot r^8=3\cdot 2^8=768"